OpenStudy (anonymous):

Can someone please help asap? I'm lost on this question I think the answer is D, am I correct? which of the following is equivalent to e+f/a-d = b/a A. a(e+f) = b(a-d) B. a+f/a-d = b/a C. ae+f = ab - d D. b(e+f) = a(a-d)

OpenStudy (anonymous):

Please someone help

OpenStudy (jdoe0001):

\(\bf \cfrac{e+f}{a-d}=\cfrac{b}{a}\qquad \textit{cross-multiplying, what would you get?}\)

OpenStudy (gudden):

Avonn ... try it once again...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@gudden are you there?

OpenStudy (gudden):

yup i am.... so try solving the whole question here in the comments and lets find it out that where do u make a mistake..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first you have to cross multiply, right?

OpenStudy (gudden):

yeah...u r .correct..till there

OpenStudy (anonymous):

a+f / a-d = b/a?

OpenStudy (gudden):

@Avonniculleyy .. u are typing the wrong equation... the correct equation ( as mentioned by u in the question ) is (e+f)/(a-d) = b/a

OpenStudy (anonymous):

Okay, what I typed was what I got as an answer

OpenStudy (gudden):

so.. what do u say about the answer now..?? which option is correct avon ??

OpenStudy (anonymous):

@gudden B. a+f/a-d = b/a

OpenStudy (gudden):

NO..... where does the " e " term go ??

OpenStudy (gudden):

how did u eliminate it..??

OpenStudy (gudden):

im now gonna solve the whole question n then we would discuss it out that where are u doing the mistake ... okay ?

OpenStudy (gudden):



OpenStudy (gudden):

so the correct option is A..